### v3

# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

Abstract.
The Model Context Protocol (MCP) is an emerging open standard that defines a unified, bi-directional communication and dynamic discovery protocol between AI models and external tools or resources, aiming to enhance interoperability and reduce fragmentation across diverse systems. This paper presents a systematic study of MCP from both architectural and security perspectives. We first define the full lifecycle of an MCP server, comprising four phases (creation, deployment, operation, and maintenance), further decomposed into 16 key activities that capture its functional evolution. Building on this lifecycle analysis, we construct a comprehensive threat taxonomy that categorizes security and privacy risks across four major attacker types: malicious developers, external attackers, malicious users, and security flaws, encompassing 16 distinct threat scenarios. To validate these risks, we develop and analyze real-world case studies that demonstrate concrete attack surfaces and vulnerability manifestations within MCP implementations. Based on these findings, the paper proposes a set of fine-grained, actionable security safeguards tailored to each lifecycle phase and threat category, offering practical guidance for secure MCP adoption. We also analyze the current MCP landscape, covering industry adoption, integration patterns, and supporting tools, to identify its technological strengths as well as existing limitations that constrain broader deployment. Finally, we outline future research and development directions aimed at strengthening MCP’s standardization, trust boundaries, and sustainable growth within the evolving ecosystem of tool-augmented AI systems. All collected data and implementation examples are publicly available at https://github.com/security-pride/MCP_Landscape.
Model Context Protocol, MCP, Vision paper, Security
$c$s: General and reference Surveys and overviewsccs: Security and privacy Software and application securityccs: Computing methodologies Artificial intelligence
1. Introduction
In recent years, the vision of autonomous AI agents capable of interacting with a wide range of tools and data sources has gained significant momentum. This progress accelerated in 2023 with the introduction of function calling by OpenAI, which allowed language models to invoke external APIs in a structured way (OpenAI, 2023b). This advancement expanded the capabilities of LLMs, enabling them to retrieve real-time data, perform computations, and interact with external systems. As function calling gained adoption, an ecosystem formed around it. OpenAI introduced the ChatGPT plugin (OpenAI, 2023a), allowing developers to build callable tools for ChatGPT. LLM app stores such as Coze (ByteDance, 2024) and Yuanqi (Tencent, 2024) have launched their plugin stores, supporting tools specifically designed for their platforms. Frameworks like LangChain (LangChain, 2022) and LlamaIndex (Liu, 2022) provided standardized tool interfaces, making it easier to integrate LLMs with external services. Other AI providers, including Anthropic, Google, and Meta, introduced similar mechanisms, further driving adoption. Despite these advancements, integrating tools remains fragmented. Developers must manually define interfaces, manage authentication, and handle execution logic for each service. Function calling mechanisms vary across platforms, requiring redundant implementations. Moreover, existing agent frameworks already support a degree of autonomous tool selection, but they generally operate over predefined or hardcoded integrations, limiting interoperability and long-term maintainability.
In late 2024, Anthropic launched the Model Context Protocol (MCP) (Anthropic, 2024b), a universal protocol for standardizing the definition, discovery, and invocation of external tools for AI applications. Drawing inspiration from the Language Server Protocol (LSP) (Gunasinghe and Marcus, 2021), MCP introduces several innovations that extend beyond conventional function calling. First, it provides a protocol-based standard that decouples tool implementation from usage, enabling developers to publish and describe external functions dynamically, independent of any single model or agent framework. Second, MCP supports dynamic discovery and schema negotiation: the client can list available tools at runtime, retrieve their capabilities, and invoke them in a uniform manner, without requiring prior hardcoding or platform-specific adapters. Third, MCP enables bi-directional communication channels, allowing not only model-to-tool requests but also tool-initiated events and notifications back to the host. Finally, MCP designs access control and capability negotiation as first-class features, offering a foundation for more auditable and secure AI-to-tool interactions. These architectural differences position MCP as more than just an advanced function-calling mechanism. It shifts the paradigm from tool bindings hardcoded per application toward an interoperable ecosystem of composable, discoverable network services. Since its release, MCP has rapidly grown into a foundational architecture for AI-native applications: thousands of independently developed MCP servers expose model-accessible interfaces to services such as GitHub (Protocol, 2024a), Slack (Protocol, 2024b), and Blender (ahujasid, 2025), while platforms like Cursor (Cursor, 2025) and Claude Desktop (Anthropic, 2024a) demonstrate how MCP-enabled clients can flexibly extend functionality by connecting to new servers on demand. This approach transforms developer tools, productivity platforms, and creative environments into a truly interoperable and multimodal AI ecosystem.
Despite the rapid adoption of MCP, its ecosystem is still in the early stages, with key areas such as security, tool discoverability, and remote deployment lacking comprehensive solutions. These issues present untapped opportunities for further research and development. Although MCP is widely recognized for its potential in the industry, it has not yet been extensively analyzed in academic research. This gap motivates our work, which provides the first in-depth analysis of the MCP ecosystem by examining its architecture and workflow, formally defining the lifecycle of MCP servers across four phases and 16 key activities, and systematically analyzing security threats from multiple attacker perspectives. Our threat taxonomy identifies four major attacker types, including malicious developers, external attackers, malicious users, and security flaws, and covers 16 representative threat scenarios. These risks, such as tool poisoning, installer spoofing, and unauthorized access, are empirically validated through real-world case studies. Through this study, we present a thorough exploration of MCP’s current landscape and offer a forward-looking vision that highlights key implications, outlines future research directions, and addresses the challenges that must be overcome to ensure its sustainable growth.
Our contributions are as follows:
  • We provide the first analysis of the MCP ecosystem, detailing its architecture, components, and workflow.
  • We identify the key components of MCP servers and define their lifecycle, encompassing the stages of creation, deployment, operation, and maintenance, across 16 key activities.
  • We construct the systematic threat taxonomy for MCP, identifying four attacker archetypes (i.e., malicious developers, external attackers, malicious users, and security flaws) and 16 threat scenarios that together reveal the MCP’s primary security exposure points.
  • We examine the current MCP ecosystem landscape, analyzing the adoption, diversity, and use cases across various industries and platforms.
  • We discuss the implications of MCP’s rapid adoption, identifying key challenges for stakeholders, and outline future research directions on security, scalability, and governance to ensure its sustainable growth.
The remainder of this paper is structured as follows: § 2 compares tool invocation with and without MCP, highlighting the motivation for this study. § 3 outlines the architecture of MCP, detailing the roles of the MCP host, client, and server, as well as the lifecycle of the MCP server. § 4 examines the current MCP landscape, focusing on key industry players and adoption trends. § 5 analyzes security and privacy risks of the MCP server and proposes mitigation strategies. § 6 explores implications, future challenges, and recommendations to enhance MCP’s scalability and security in dynamic AI environments. § 7 reviews prior work on tool integration and security in LLM applications. Finally, § 8 concludes the whole paper.
2. Background and Motivation
2.1. AI Tooling
Before the introduction of MCP, AI applications relied on various methods, such as manual API wiring, plugin-based interfaces, and agent frameworks, to interact with external tools. As shown in Figure 1, these approaches required integrating each external service with a specific API, leading to increased complexity and limited scalability. MCP addresses these challenges by providing a standardized protocol that enables seamless and flexible interaction with multiple tools.
Figure 1. Tool invocation with and without MCP. Without MCP, an AI application interacts with external tools and resources such as web services, databases, and local files through specific APIs. With MCP, the AI application functions as an MCP client that communicates with an MCP server using the MCP protocol, which provides a unified interface for tool access.
2.1.1. Manual API Wiring
In traditional implementations, developers had to establish manual API connections for each tool or service that an AI application interacted with. This process required custom authentication, data transformation, and error handling for every integration. As the number of APIs increased, the maintenance burden became significant, often leading to tightly coupled and fragile systems that were difficult to scale or modify. MCP eliminates this complexity by offering a unified interface, allowing AI models to connect with multiple tools dynamically without the need for custom API wiring.
2.1.2. Standardized Plugin Interfaces
To reduce the complexity of manual wiring, plugin-based interfaces such as OpenAI ChatGPT Plugins, introduced in November 2023 (OpenAI, 2023a), allowed AI models to connect with external tools through standardized API schemas like OpenAPI. For example, in the OpenAI Plugin ecosystem, plugins like Zapier allowed models to perform predefined actions, such as sending emails or updating CRM records. However, these interactions were often one-directional and could not maintain state or coordinate multiple steps in a task. There are also new LLM app stores (Zhao et al., 2024) that offer a web services plugin store, like ByteDance Coze (ByteDance, 2024) and Tencent Yuanqi (Tencent, 2024). Although these platforms increased the number of tools available, they also produced separate ecosystems where plugins are platform-specific, which restricts limiting cross-platform compatibility and requiring duplicate maintenance efforts. MCP stands out by being open-source and platform-agnostic, supporting bi-directional communication channels that allow not only model-to-tool invocations but also tool-initiated events and notifications, which are not possible in one-directional plugin designs.
2.1.3. AI Agent Tool Integration
The emergence of AI agent frameworks like LangChain (LangChain, 2022) and similar tool orchestration frameworks provided a structured way for models to invoke external tools through predefined interfaces, improving automation and adaptability (Xi et al., 2023). However, integrating and maintaining these tools remained largely manual, requiring custom implementations and increasing complexity as the number of tools grew. MCP simplifies this process by introducing a protocol-level abstraction that unifies how tools are described and discovered across platforms. This protocol, in contrast to framework-specific integrations, makes it possible for tools created by various developers to work together, eliminating the need for duplicate maintenance and promoting a shared ecosystem.
2.1.4. Retrieval-Augmented Generation (RAG) and Vector Database.
Contextual information retrieval methods, such as RAG, leverage vector-based search to retrieve relevant knowledge from databases or knowledge bases, enabling models to supplement responses with up-to-date information (Fan et al., 2024; Cuconasu et al., 2024). While this approach addressed the problem of knowledge cutoff and improved model accuracy, it was limited to passive retrieval of information. It did not inherently allow models to perform active operations, such as modifying data. For example, a RAG-based system could retrieve relevant sections from a product documentation database to assist a customer support AI. However, if the AI needed to update customer records or escalate an issue to human support, it can’t take action beyond providing textual responses. MCP extends beyond passive information retrieval by providing a standardized protocol for both retrieval and action, allowing AI systems to combine knowledge access with secure tool execution under a unified framework.
2.2. Motivation
MCP has rapidly gained traction in the AI community due to its ability to standardize how AI models interact with external tools, fetch data, and execute operations. By addressing the limitations of manual API wiring, plugin interfaces, and agent frameworks, MCP redefines AI-to-tool interactions by enabling interoperable, secure, and maintainable workflows across heterogeneous environments. Unlike prior approaches, MCP incorporates access control, capability negotiation, and schema discovery as protocol primitives, which distinguishes it from conventional function-calling mechanisms. Despite its growing adoption and promising potential, MCP is still in its early stages, with an evolving ecosystem that remains incomplete. Many key aspects, such as security and tool discoverability, are yet to be fully addressed, leaving ample room for future research and improvement. Moreover, while MCP has gained rapid adoption in the industry, it is still largely unexplored in academia.
Motivated by this gap, this paper is the first to analyze the current MCP landscape, examine its emerging ecosystem, and identify potential security risks. Additionally, we outline a vision for MCP’s future development and highlight the key challenges that must be addressed to support its long-term success.
3. MCP Architecture
3.1. Core Components
The MCP architecture is composed of three core components: MCP host, MCP client, and MCP server. These components collaborate to facilitate seamless communication between AI applications, external tools, and data sources, ensuring that operations are secure and properly managed. As shown in Figure 2, in a typical workflow, the user sends a prompt to the MCP client, which analyzes the intent, selects the appropriate tools via the MCP server, and invokes external APIs to retrieve and process the required information before notifying the user of the results.
Figure 2. The workflow of MCP. A user prompt is processed through a series of stages involving intent analysis, tool selection, and API invocation across the MCP host, client, and server. The MCP server provides tools, resources, and prompts that enable interaction with external data sources such as web services, databases, and local files. The notation “1:1” in the transfer layer indicates a one-to-one communication link between each MCP client and MCP server during request and response exchange.
3.1.1. MCP Host
The MCP host is an AI application that provides the environment for executing AI-based tasks while running the MCP client. It integrates interactive tools and data to enable smooth communication with external services. Examples include Claude Desktop for AI-assisted content creation, Cursor, an AI-powered IDE for code completion and software development, and AI agents that function as autonomous systems for executing complex tasks. The MCP host hosts the MCP client and ensures communication with external MCP servers.
3.1.2. MCP Client
In the MCP host–client–server architecture, the MCP client acts as an intermediary component within the host environment, maintaining a one-to-one communication link with its corresponding MCP server. Operating on behalf of the host, the client initiates requests to the MCP server, queries available functions, and retrieves responses that describe the server’s tools, resources, and capabilities. This design enables the host to seamlessly access and utilize external functionalities provided by MCP servers. In addition to managing requests and responses, the MCP client processes notifications from MCP servers, providing real-time updates about task progress and system status. It also performs sampling to gather data on tool usage and performance, enabling optimization and informed decision-making. The MCP client communicates through the transport layer with MCP servers, facilitating secure, reliable data exchange and smooth interaction between the host and external resources.
3.1.3. MCP Server
The MCP server enables the MCP host and client to access external systems and execute operations, offering three core capabilities: tools, resources, and prompts.
  • Tools: Enabling external operations. Tools in MCP enable the server to invoke external services and APIs to execute operations on behalf of AI models. When a host application (such as an AI assistant) needs to perform an operation, it first queries the MCP server through the client to obtain the list of available tools and their capabilities. Based on the task context, the host then selects an appropriate tool and issues an invocation request via the client. The MCP server executes the corresponding operation through the external service or API and returns the result to the client, which forwards it to the host. For example, if an AI model requires real-time weather data or sentiment analysis, the host identifies the corresponding tool from the server’s advertised capabilities, the server performs the API call, and the resulting data is delivered back to the host. Unlike traditional function-calling interfaces that are confined within a single model or framework, tools in MCP are described and accessed through a standardized, model-agnostic protocol. This design allows tools to be dynamically discovered, described, and invoked across heterogeneous AI systems, ensuring cross-platform interoperability rather than provider-specific integration. Furthermore, MCP tools support bi-directional communication between the host and the service, enabling richer interactions such as asynchronous updates or event streaming. Once configured, these tools adhere to a supply-and-consume model that promotes modularity and reusability, allowing independently developed tools to interoperate seamlessly and improving system efficiency and extensibility.
  • Resources: Exposing data to AI models. Resources provide access to structured and unstructured datasets that the MCP server can expose to AI models. These datasets may come from local storage, databases, or cloud platforms. When an AI model requests specific data, the MCP server retrieves and processes the relevant information, enabling the model to make data-driven decisions. For example, a recommendation system may access customer interaction logs, or a document summarization task may query a text repository.
  • Prompts: Reusable templates for workflow optimization. Prompts are predefined templates and workflows that the MCP server generates and maintains to optimize AI responses and streamline repetitive tasks. They ensure consistency in responses and improve task execution efficiency. For instance, a customer support chatbot may use prompt templates to provide uniform and accurate responses, while an annotation task may rely on predefined prompts to maintain consistency in data labeling.
3.2. Transport Layer and Communication
The transport layer ensures secure, bidirectional communication, allowing for real-time interaction and efficient data exchange between the host environment and external systems. The transport layer manages the transmission of initial requests from the client, the delivery of server responses detailing available capabilities, and the exchange of notifications that keep the client informed of ongoing updates. Communication between the MCP client and the MCP server follows a structured process, beginning with an initial request from the client to query the server’s functionalities. Upon receiving the request, the server responds with an initial response listing the available tools, resources, and prompts the client can leverage. Once the connection is established, the system maintains a continuous exchange of notifications to ensure that changes in server status or updates are communicated back to the client in real time. This structured communication ensures high-performance interactions and keeps AI models synchronized with external resources, enhancing the effectiveness of AI applications.
3.3. MCP Server Lifecycle
Figure 3 summarizes the complete lifecycle of an MCP server from the server’s perspective. This lifecycle is primarily derived from the official protocol documentation and a systematic analysis of actual MCP operational workflows. Based on the transitions of primary participant roles, the lifecycle is divided into four sequential phases: creation, deployment, operation, and maintenance. In the creation phase, the main actor is the developer, who defines metadata, declares capabilities, and implements the MCP server. The deployment phase covers the process in which the developer releases the server to a public platform, and users deploy it to an MCP host while the client establishes its connection. The operation phase corresponds to the runtime period when users actively interact with the server through the MCP system. Finally, the maintenance phase involves version iteration, configuration updates, and continuous optimization of the deployed MCP server. The following subsection introduces the key activities of each phase in detail.
Figure 3. MCP server components and lifecycle. The upper part presents the time-ordered interaction flow among the developer, MCP server, host, client, user, and external resources. The lower part summarizes the main lifecycle phases, including creation, deployment, operation, and maintenance. The numbered circles (e.g., ①②③) indicate sequential actions in the upper process, which correspond to key phases in the lifecycle.
3.3.1. MCP Server Components
The MCP server is responsible for managing external tools, data sources, and workflows, providing AI models with the necessary resources to perform tasks efficiently and securely. It comprises several key components that ensure smooth and effective operations. Metadata includes essential information about the server, such as its name, version, and description, allowing clients to identify and interact with the appropriate server. Configuration involves the source code, configuration files, and manifest, which define the server’s operational parameters, environment settings, and security policies. Tool list stores a catalog of available tools, detailing their functionalities, input-output formats, and access permissions, ensuring proper tool management and security. Resources list governs access to external data sources, including web APIs, databases, and local files, specifying allowed endpoints and their associated permissions. Finally, Prompt templates include pre-configured task templates and workflows that enhance the efficiency of AI models in executing complex operations. These components enable MCP servers to provide seamless tool integration, data retrieval, and task orchestration for AI-powered applications.
3.3.2. Creation Phase
The creation phase is the starting point of the MCP server lifecycle. During this stage, developers establish the server’s essential data structures and capabilities, transforming conceptual requirements into a fully defined and executable MCP component. The result is a standardized service that can be reliably integrated into subsequent deployment and operation processes. Metadata definition focuses on describing the server’s identity through its name, version, description, and supported protocol version. These data form the foundation for interoperability and version control. Capability declaration specifies the standardized functions the server provides, such as tools, resources, or prompts, together with their operational boundaries and permission requirements. This step determines how the server will respond to standardized requests. Vague or inaccurate declarations may lead to capability misuse or potential security violations. Many security vulnerabilities identified in MCP deployments originate from flaws introduced at this stage. Therefore, scanning and validation of capability declarations constitute an essential step in strengthening MCP server security. Code implementation connects the declared capabilities with concrete request handlers that process input data, execute corresponding logic, and generate structured results. The quality and security of this implementation directly influence the stability and safety of all exposed functions. Poorly designed or insecure code can introduce vulnerabilities that compromise both reliability and user trust. Slash command definition establishes optional user‑interaction commands that correspond to specific prompts. Clear and coherent command definitions enhance usability and improve the flexibility and intuitiveness of prompt invocation during subsequent operations.
3.3.3. Deployment Phase
In this stage, the server that has been designed and implemented is prepared, packaged, and released into an operational environment where clients and hosts can interact with it through standardized interfaces. MCP server release involves packaging the finalized server codebase, configuration files, and metadata into a distributable form. This step may include version tagging, dependency documentation, and integrity verification to guarantee that the deployed artifact remains consistent with the validated build. Developers can subsequently publish their packaged servers to various MCP server markets, as illustrated in Table 2, allowing end users to discover and install servers directly from these repositories. Most existing markets are maintained by third‑party platforms, but Anthropic has also begun developing an official MCP registry aimed at providing verified listings, enhanced security trust, and unified distribution management for the MCP ecosystem. Installer deployment handles the distribution and installation of the server package within target systems. Installers can use container images, package managers, or automated scripts to streamline deployment. Clear installation workflows reduce setup errors and make the integration process predictable. Security validation of installers, such as checksum verification and signature authentication, is necessary to prevent tampering or injection of malicious binaries. Environment setup ensures that the deployed MCP server operates under the correct runtime configuration. This includes defining environment variables, access credentials, logging policies, and network permissions required for communication with clients and hosts. Configuration isolation and principle‑of‑least‑privilege practices help mitigate risks of unauthorized access or data leakage during runtime. Tool registration finalizes deployment by registering the server’s capabilities with the hosting application or orchestration system. During this process, tools, resources, and prompts declared during creation are discovered and made available to connected clients.
3.3.4. Operation Phase
The operation phase represents the runtime stage of the MCP server lifecycle, where the deployed server actively interacts with users, clients, and external resources. During this phase, the server interprets user intent, mediates resource access, invokes registered tools, and manages ongoing sessions to deliver reliable and contextual responses. Intent analysis is the initial operation stage in which user inputs are parsed, contextualized, and mapped to the most suitable MCP capabilities. The host or client forwards user requests to the MCP server, where intent interpretation logic determines whether to trigger a tool, retrieve a resource, or engage a prompt. Accurate intent analysis directly affects usability and effectiveness. Misinterpretation can lead to incorrect tool execution or unnecessary resource calls, potentially increasing latency or amplifying risk exposure. External resource access occurs when the server needs to obtain supplementary data from third‑party systems or knowledge repositories. Such access is handled via predefined resource interfaces and is strictly governed by security permissions defined during creation. Each access request must comply with authentication, authorization, and sandboxing policies to prevent data leakage or external dependency compromise. Tool invocation is the execution stage where the MCP server or client triggers a registered tool according to the interpreted intent. Each tool invocation includes structured parameter passing, execution monitoring, and result serialization before returning structured outputs to the user session. Reliability and isolation at this stage determine runtime stability. Session management maintains the logical continuity between user interactions and server processes. It includes establishing, monitoring, and closing session contexts that connect the MCP host, client, and user interface.
3.3.5. Maintenance Phase
After deployment and operation, an MCP server must be periodically maintained to ensure its reliability, security, and compliance with evolving system requirements. Version control ensures that all updates to the MCP server, including bug fixes and capability extensions, are tracked through an auditable revision system. Proper version management enables rollback to stable releases, facilitates controlled feature rollout, and supports compatibility testing across client environments. Configuration change management governs modifications to runtime parameters, environmental variables, or authorization policies. All configuration adjustments should follow a controlled workflow, such as change‑request approval and pre‑deployment validation, to prevent inadvertent disruptions. Access audit records and reviews all authentication and authorization events related to the MCP server. This includes tracking user sessions, privilege escalations, and external resource permissions utilized during operation. Regular access auditing allows administrators to identify abnormal patterns, enforce least‑privilege policies, and comply with organizational or regulatory security frameworks. Log audit focuses on continuous collection and analysis of operational logs generated by the server, tools, and clients. Centralized log aggregation and integrity‑protected storage support forensic traceability and incident response. Automated log analysis, through anomaly detection or correlation with known attack signatures, can reveal early indicators of compromise or performance degradation. Effective log auditing transforms runtime data into actionable intelligence for system governance.
4. Current Landscape
4.1. Ecosystem Overview
4.1.1. Key Adopters.
Table 1 demonstrates how MCP has gained significant traction across diverse sectors, signaling its growing importance in enabling seamless AI-to-tool interactions. The dataset summarized in this table was compiled through manual inspection, drawing on official documentation from the earliest MCP supporters and extended via community discussions and repository mining. We primarily included mature products and companies with verifiable MCP integration. Although this collection provides a representative snapshot of the current MCP landscape, we acknowledge that it is not exhaustive. To enhance transparency and facilitate future updates, the dataset will be maintained as a public repository [^1] https://github.com/security-pride/MCP_Landscape, enabling ongoing community contributions and periodic verification.
Table 1. Overview of MCP ecosystem adoption (updated to Sept. 2025).
Category Company/Product Key Features or Use Cases
AI Models and Frameworks Anthropic (Claude) (Anthropic, 2024a) Full MCP support in the desktop version, enabling interaction with external tools.
OpenAI (OpenAI, 2025b) MCP support integrated across products and within the Agent SDK for seamless interoperability.
Google DeepMind (DeepMind, 2025) Added MCP protocol support for the Gemini model family, enabling standardized tool invocation.
Baidu Maps (Maps, 2025) API integration using MCP to access geolocation and spatial services.
Blender MCP (MCP, 2025b) Enables Blender and Unity 3D model generation via natural language commands.
Developer Tools Replit (Replit, 2025) AI-assisted development environment with integrated MCP tools.
Microsoft Copilot Studio (Studio, 2025) Officially announced MCP support in March 2025, enabling unified tool integration.
Sourcegraph Cody (Cody, 2025) Implements MCP through OpenCTX for resource integration.
Codeium (Codeium, 2025) Adds MCP support for coding assistants to facilitate cross-system tasks.
Cursor (Cursor, 2025) MCP tool integration in Cursor Composer for seamless code execution.
Cline (Cline, 2025) VS Code coding agent that manages MCP tools and servers.
IDEs/Editors Zed (Zed, 2025) Provides slash commands and tool integration based on MCP.
JetBrains (JetBrains, 2025) Integrates MCP for IDE-based AI tooling.
Windsurf Editor (Editor, 2025) AI-assisted IDE with MCP tool interaction.
TheiaAI/TheiaIDE (TheiaAI/TheiaIDE, 2025) Enables MCP server interaction for AI-powered tools.
Emacs MCP (MCP, 2025a) Enhances AI functionality in Emacs by supporting MCP tool invocation.
OpenSumi (OpenSumi, 2025) Supports MCP tools in IDEs and enables seamless AI tool integration.
Cloud Platforms and Services Cloudflare (Cloudflare, 2025) Provides remote MCP server hosting, OAuth integration, and scalable multi-tenant infrastructure.
Tencent Cloud (Cloud, 2025c) Added MCP plugin and transport support, enabling AI SDK-level integration.
Alibaba Cloud Bailian (Cloud, 2025a) Introduced full lifecycle MCP service for large-scale batch integration.
Huawei Cloud (Cloud, 2025b) Launched AI-native run platform with built-in MCP module to accelerate trusted AI deployment.
Block (Square) ((2025), Square) Uses MCP to enhance data processing efficiency for financial platforms.
Stripe (Stripe, 2025) Exposes payment APIs via MCP for seamless AI integration.
Alipay / Ant Group (Alipay, 2025; Group, 2025) Released “Alipay MCP Server” and “MCP Zone”, offering unified payment and tool orchestration.
Web Automation and Data Apify MCP Tester (Tester, 2025) Connects to any MCP server using SSE for API testing.
LibreChat (LibreChat, 2025) Extends the current tool ecosystem through MCP integration.
Baidu Create Conference (Baidu, 2025) Established a dedicated MCP ecosystem forum to promote developer collaboration.
Notably, leading AI companies such as Anthropic (Anthropic, 2024a) and OpenAI (OpenAI, 2025b) have integrated MCP to enhance agent capabilities and improve multi-step task execution. This adoption by industry pioneers has set a precedent, encouraging other major players to follow suit. Chinese tech giants like Baidu (Maps, 2025) have also incorporated MCP into their ecosystems, highlighting the protocol’s potential to standardize AI workflows across global markets. Developer tools and IDEs, including Replit (Replit, 2025), Microsoft Copilot Studio (Studio, 2025), JetBrains (JetBrains, 2025), and TheiaIDE (TheiaAI/TheiaIDE, 2025), leverage MCP to facilitate agentic workflows and streamline cross-platform operations. This trend indicates a shift toward embedding MCP in developer environments to enhance productivity and reduce manual integration efforts. Furthermore, cloud platforms like Cloudflare (Cloudflare, 2025) and financial service providers such as Block (Square) ((2025), Square) and Stripe (Stripe, 2025) are exploring MCP to improve security, scalability, and governance in multi-tenant environments. The widespread adoption of MCP by these industry leaders not only highlights its growing relevance but also points to its potential as a foundational layer in AI-powered ecosystems. As more companies integrate MCP into their operations, the protocol is set to play a central role in shaping the future of AI tool integration. Looking ahead, MCP is poised to become a key enabler of AI-driven workflows, driving more secure, scalable, and efficient AI ecosystems across industries.
While these examples show clear momentum, the overall MCP ecosystem remains uneven in maturity. Many community-hosted servers are small experiments or early prototypes. Only a limited number demonstrate strong reliability, stable maintenance, or standardized documentation. A deeper issue lies in platform governance. Many large internet platforms operate as walled gardens. They treat data as a controlled resource and limit access to interfaces or user behavior data to maintain a competitive advantage. In such systems, the open and composable design that MCP promotes may face natural resistance. We note that platforms might only expose low-risk or secondary functions through MCP connections. For example, an online map service might allow MCP access to public location data or route information, but not to personalized navigation history or user-generated content. Similarly, a social media platform could open an MCP interface for basic profile lookup or public post retrieval, while keeping private messages and recommendation algorithms restricted. These partial integrations show that platform openness often depends on strategic and privacy considerations rather than technical feasibility. This reflects a tension between open interoperability and proprietary data control. The success of MCP will partly depend on whether large platforms are willing to share meaningful data and support cross-system collaboration.
4.1.2. Community-Driven MCP Servers.
Although Anthropic has not yet released an official MCP marketplace, the broader community has actively filled this gap by creating numerous independent MCP server collections and discovery platforms. As summarized in Table 2, we identified and catalogued all publicly available MCP server directories accessible as of September 2025. These include a diverse range of deployment mode, including websites, GitHub repositories, and desktop applications, demonstrating how third-party developers are collectively working to establish a complete and sustainable MCP ecosystem. The data presented in Table 2 were collected and verified manually. We began by surveying known MCP-related repositories on GitHub and popular community forums, followed by direct examination of listed directory or application to confirm its activity, accessibility, and declared number of hosted servers. Most platforms explicitly report total server counts, which we cross-checked against their public listings; where such data were missing, we manually counts from available catalog pages. This collection process yielded a consolidated dataset encompassing 26 major MCP collections. Platforms such as MCP.so (mcp.so, 2025), Glama (glama.ai, 2025), and PulseMCP (et al., 2025) host thousands of servers, allowing users to discover and integrate a wide range of tools and services. These community-driven platforms have significantly accelerated the adoption of MCP by providing accessible repositories where developers can publish, manage, and share their MCP servers. Desktop-based solutions like Dockmaster (mcp dockmater, 2025) and Toolbase (gching, 2025) further enhance local MCP deployment capabilities, empowering developers to manage and experiment with servers in isolated environments.
Despite this rapid growth, the quality of community-maintained MCP markets remains uneven. Platforms like MCP.so host thousands of entries but lack formal security or identity verification mechanisms. To assess reliability, we randomly sampled 300 servers listed on MCP.so. Among them, 30 contained the term “MCP” in the project title but did not refer to the Model Context Protocol, and 18 were in active development or unavailable at the time of inspection. These findings suggest that large community directories may overstate their effective numbers, and server quality varies widely across listings. In contrast, platforms such as mcp-get implement verification and signing processes, but the number of verified servers remains very small and user adoption is limited. The overall landscape therefore reflects a trade‑off between openness and quality control: while community initiatives accelerate ecosystem growth, they also highlight the need for stronger validation standards and sustainable curation practices.
Table 2. Overview of MCP server collections and deployment modes (As of Sept. 2025).
Collection Author Mode # Servers URL
MCPWorld Baidu Website 26,404 mcpworld.com
MCP.so mcpso Website 16,592 mcp.so
MCP Servers Repository mcprepository Website 13,596 mcprepository.com
AIbase MCP AIbase Website 12,448 mcp.aibase.com
Glama glama.ai Website 9,415 glama.ai
Smithery Henry Mao Website 6,888 smithery.ai
PulseMCP Antanavicius et al. Website 6,072 pulsemcp.com
ModelScope MCP Marketplace modelscope Website 5,441 modelscope.cn/mcp
Awesome MCP Servers wong2 Website 2,402 mcpservers.org
Cursor Directory MCP Cursor Website 1,800 cursor.directory/mcp
Official Collection1 Anthropic GitHub Repo 1,204 modelcontextprotocol/servers
AiMCP Hekmon Website 907 aimcp.info
Dockmaster mcp-dockmaster Desktop App 516 mcp-dockmaster.com
MCP Market mcpmarket.com Website 463 mcpmarket.com
MCP.run mcp.run Website 242 mcp.run
Awesome MCP Servers Stephen Akinyemi GitHub Repo 217 appcypher/mcp-servers
CLine MCP Marketplace Cline Website 154 cline.bot/mcp-marketplace
Bailian MCP Market Aliyun Website 151 bailian.console.aliyun.com
OpenTools opentoolsteam Website 148 opentools.com
Awesome Remote MCP Servers JAW9C GitHub Repo 79 jaw9c/awesome-remote-mcp-servers
MCP Server Hub mcpserverhub Website 71 mcpserverhub.com
mcp-get Michael Latman Website 59 mcp-get.com
MCP Marketplace Higress.ai Website 50 mcp.higress.ai
Toolbase gching Desktop App 24 gettoolbase.ai
make inference mkinf Website 23 mkinf.io
Awesome Crypto MCP Servers Luke Fan GitHub Repo 12 badkk/crypto-mcp-servers
  • 1
    Official Collection refers to the list of MCP servers curated by Anthropic. Anthropic has also launched the community-driven MCP Registry, a preview service that provides a centralized directory for discovering MCP servers (https://github.com/modelcontextprotocol/registry).
4.1.3. SDKs and Tools.
With the continuous growth of community-driven tools and official SDKs, the MCP ecosystem is becoming increasingly accessible, allowing developers to integrate MCP into various applications and workflows efficiently. Official SDKs are available in multiple languages, including TypeScript, Python, Java, Kotlin, and C#, providing developers with versatile options to implement MCP in different environments. In addition to official SDKs, the community has contributed numerous frameworks and utilities that simplify MCP server development. Tools such as EasyMCP and FastMCP offer lightweight TypeScript-based solutions for quickly building MCP servers, while FastAPI to MCP Auto Generator enables the seamless exposure of FastAPI endpoints as MCP tools. For more complex scenarios, Foxy Contexts provides a Golang-based library to build MCP servers, and Higress MCP Server Hosting extends the API Gateway (based on Envoy) to host MCP servers with wasm plugins. Server generation and management platforms such as Mintlify, Speakeasy, and Stainless further enhance the ecosystem by automating MCP server generation, providing curated MCP server lists, and enabling faster deployment with minimal manual intervention. These platforms empower organizations to rapidly create and manage secure and well-documented MCP servers.
4.2. Use Cases
MCP has become a vital tool for AI applications to effectively communicate with external tools, APIs, and systems. By standardizing interactions, MCP simplifies complex workflows, boosting the efficiency of AI-driven applications. Below, we explore three key platforms (i.e., OpenAI, Cursor, and Cloudflare) that have successfully integrated MCP, highlighting their distinct use cases.
4.2.1. OpenAI: MCP Integration in AI Agents and SDKs
We chose OpenAI as the case study because it holds significant influence in the field of agent development. The Agents SDK (OpenAI, 2025a), as a mainstream development framework, covers a wide range of real-world applications. It incorporates basic capabilities such as managing conversations, invoking tools, task delegation, input and output verification, and session recording. Developers only need to write a small amount of Python code to enable the intelligent agent to handle multiple rounds of conversations automatically, flexibly call various tools, and even allow multiple intelligent agents to collaborate to complete complex tasks. After the MCP was added, the Agents SDK became more efficient in terms of tool integration. In the past, developers had to customize the connection logic separately for each new tool or external service. This was not only time-consuming but also prone to errors. Now, with the help of MCP, developers only need to configure the address of the MCP tool, and the agent can automatically discover, connect to, and call it, regardless of where these tools are deployed. The entire integration process has become more unified, and maintaining and extending new tools has become much simpler. ChatGPT currently supports MCP in developer mode. Users can directly connect MCP tools within ChatGPT, enabling them to do more than just query data. They can also write to external systems, trigger automated tasks, or link multiple tools together to perform complex operations. The introduction of MCP has transformed ChatGPT from a past question-and-answer assistant into a platform that can continuously expand its capabilities.
During the process of promoting and practicing MCP, OpenAI has accumulated rich experience and provided many applicable practices for the industry. Through this case, we can observe how MCP helps developers avoid detours and accelerate the transformation of ideas into practical applications. This not only makes the integration of tools smoother but also makes AI systems more open and flexible.
4.2.2. Cursor: Enhancing Software Development with MCP-Powered Code Assistants
In recent years, AI-assisted programming tools have significantly simplified the software development process. Tools like Cursor, Claude Code, and Cline enable developers to interact with the system using natural language. These assistants can automatically generate code, assist with debugging, and also perform structural reconfiguration. As a result, developers can focus more on business logic, leading to a significant improvement in development efficiency and code quality. The integration of the MCP protocol has further expanded the capabilities of AI-assisted programming tools. Taking Cursor as an example, MCP enables it to directly access external APIs, code repositories, and various automation tools. In practical applications, this process typically includes steps such as instruction parsing, tool scheduling, task distribution, and result feedback. Developers simply input tasks in Cursor, such as “Help me perform a page visit and take a screenshot on qq.com”, and Cursor will analyze whether external tool support is needed and request services from the backend through the MCP protocol. The MCP server is responsible for scheduling the appropriate tools, such as Playwright MCP, to automatically complete tasks like page visits and screenshots, and return the results to the agent, ultimately presenting them to the developer. Even for more complex automated testing tasks, such as form submissions, AJAX requests, or multi-page navigation, developers do not need to leave the IDE environment. All processes can be automated and standardized.
The introduction of MCP has greatly enhanced the flexibility and scalability of Cursor. Developers no longer need to write repetitive adaptation codes for each new tool; they only need to declare the tool addresses to access different types of services, regardless of whether these tools are running locally, remotely, or in the cloud. This mechanism reduces the difficulty of expansion and maintenance. Cursor can also automatically coordinate different tools based on the current development scenario, helping developers complete multiple operations. As a result, the development experience becomes smoother and the efficiency is significantly improved. Of course, the integration of MCP also brings some challenges. For example, the data compatibility between different tools needs to be addressed, and permission management and security isolation are even more important. Distributed calls may cause network delays, which have an impact on the overall experience. Despite these issues, after integrating MCP, the intelligence and openness of the development environment have significantly improved. This case helps us understand the profound impact of MCP on actual software development and provides a reference for the future development of the industry.
4.2.3. Cloudflare: Remote MCP Server Hosting and Scalability
Cloudflare transforms MCP from a local-only technology into a scalable, cloud-based solution, as illustrated in Figure 4. By hosting MCP servers in the cloud, Cloudflare removes the need for users to configure and maintain servers on their own machines. This shift lowers the technical barrier for both developers and end users. Cloudflare integrates managed authentication with OAuth 2.0, which ensures that only authorized agents and users can access MCP servers and the tools they provide. The platform also supports persistent state and secure data storage through technologies like Durable Objects and Workers KV. Each MCP session can reliably maintain its own data, even as usage grows or as users switch devices. With these capabilities, MCP servers running on Cloudflare can connect to external APIs, automate multi-step workflows, and interact with a broad range of third-party services. Developers can focus on building useful features, while Cloudflare handles critical concerns such as security, scaling, and network connectivity. This multi-tenant design allows different users or organizations to safely share the same infrastructure without the risk of data leakage.
Figure 4. Architecture of local and remote MCP communication via proxy.
Deploying MCP servers on a cloud platform introduces certain challenges. Developers must carefully define authorization scopes to limit access to sensitive resources. They also need to address privacy and data residency requirements for users in different regions. Real-time agent workflows may require special attention to maintain low latency and stable connections. Cloudflare provides tools to address many of these concerns, but secure tool design and session management remain important responsibilities for developers. Moving MCP to the cloud with Cloudflare extends the value of the protocol to a much wider audience. Both technical and non-technical users can now access secure, scalable AI-powered tools from any device. This cloud-native approach helps MCP become a universal interface for safe, extensible, and cross-device AI automation. Cloudflare’s platform shows how modern cloud infrastructure can make advanced AI agent ecosystems accessible and practical in real-world applications.
The adoption of MCP by platforms like OpenAI, Cursor, and Cloudflare highlights its flexibility and growing role in AI-driven workflows, enhancing efficiency, adaptability, and scalability across development tools, enterprise applications, and cloud services.
5. Security and Privacy Analysis
This section provides a comprehensive analysis of the security and privacy risks in the MCP ecosystem. To systematically describe potential vulnerabilities, we categorize the threats based on the type of attacker, the origin of the threat, and the possible attack consequences. As summarized in Table 3, our extended threat taxonomy covers four major attacker types: malicious developers, external attackers, malicious users, and general security flaws, each mapped to specific threat origins that may occur or manifest across multiple stages of the MCP lifecycle. Rather than rigidly assigning risks to a single phase such as creation, deployment, operation, or maintenance, we emphasize analyzing their origins, since attacks like tool poisoning and rug pulls are introduced at the creation stage but triggered at operation. This perspective helps us reason about mitigation strategies from the source, enabling more effective control and prevention of security incidents.
To further reveal the security vulnerabilities of the MCP ecosystem and validate our analysis, we constructed proof-of-concept (PoC) MCP servers corresponding to each identified risk type within an isolated environment. We also implemented a custom MCP host based on the official MCP SDK to establish controlled connections with these servers. Since the goal of this PoC is to demonstrate security risks and feasibility rather than to evaluate attack success rates or the behavioral differences among various base LLMs, we leave such comprehensive evaluations for future work.
Table 3. Threats, origins, and consequences across different attacker types.
Type Security risk Section Threat Origin Attack Consequence
Malicious
Developer
Namespace Typosquatting § 5.1.1 (1) Metadata Definition Installation of malicious server, supply chain compromise
Tool Name Conflict § 5.1.2 (1) Capability Declaration Ambiguity, wrong tool execution, privilege escalation
Preference Manipulation § 5.1.3 (1) Capability Declaration Unsafe defaults exploited, misuse of features
Tool Poisoning § 5.1.4 (1) Capability Declaration Hidden malicious payload executed
Rug Pulls § 5.1.5 (1) Capability Declaration Service disruption, loss of trust
Cross-Server Shadowing § 5.1.6 (1) Capability Declaration Malicious functionality hidden, lateral exploitation
Command Injection § 5.1.7 (1) Code Implementation Arbitrary command execution, system compromise
External
Attacker
Installer Spoofing § 5.2.1 (2) Installer Deployment Deployment of compromised MCP server
Indirect Prompt Injection § 5.2.2 (3) External Resource Access Malicious instructions injected into LLM workflow
Malicious
User
Credential Theft § 5.3.1 (3) Tool Invocation Unauthorized access to sensitive data and resources
Sandbox Escape § 5.3.2 (3) Tool Invocation Escape from isolation, host system compromise
Tool Chaining Abuse § 5.3.3 (3) Tool Invocation Abuse multiple tools for data exfiltration or escalation
Unauthorized Access § 5.3.4 (3) Session Management Session hijacking, impersonation of legitimate users
Security
Flaws
Vulnerable Versions § 5.4.1 (4) Version Control Exploitation of known CVEs, remote code execution
Privilege Persistence § 5.4.2 (4) Version Control Retained unauthorized rights, long-term compromise
Configuration Drift § 5.4.3 (4) Configuration Change Misconfiguration exposes sensitive services
  • 1
    Numbers in parentheses indicate the threat origin stage: (1) Creation, (2) Deployment, (3) Operation, and (4) Maintenance.
5.1. Malicious Developer
5.1.1. Namespace Typosquatting
Namespace typosquatting refers to a form of server name collision in which a malicious actor registers an MCP server with a name that is identical or deceptively similar to that of a legitimate one, tricking users or host applications into selecting the malicious server during deployment or runtime. During the deployment stage, end users typically install MCP servers from the public MCP market, relying mainly on the server’s name and description. During the runtime stage, the MCP host selects from available servers advertised by clients, again based primarily on those textual identifiers. This process makes both users and hosts susceptible to impersonation attacks. Once a compromised server is installed, it can mislead AI agents and clients into invoking the malicious instance, potentially exposing sensitive data, executing unauthorized commands, or disrupting workflows.
While typosquatting is a known issue in ecosystems such as package managers (e.g., npm, PyPI), plugin systems (e.g., VS Code, Chrome extensions), and cloud frameworks, it manifests differently in the MCP context. In traditional systems, users install packages through explicit manual choices, with human oversight. In MCP, interactions are often automated and mediated by AI agents or orchestration hosts. Server selection can occur dynamically at runtime, based solely on identifiers in prompts or capability metadata. This automation amplifies the potential harm: a single misselection can propagate incorrect behaviors or data leaks across multiple downstream chains of AI calls. To illustrate this risk, we constructed two MCP servers with visually similar names, github-mcp and mcp-github, as shown in Figure 5. In our demonstration, the AI application selected between the available servers solely based on names and descriptions, showing no preference or verification mechanism. Once the malicious mcp-github server was chosen, it was able to intercept commit operations and exfiltrate authentication tokens and repository data while appearing to function normally. This example demonstrates how identity ambiguity in MCP naming can lead to silent compromise, and why stronger namespace validation and signing mechanisms are required for secure interoperability.
Figure 5. Example of namespace typosquatting. A legitimate MCP server named github‑mcp is impersonated by a malicious variant mcp‑github. When users or AI applications select an MCP server, the similar names can cause confusion. If the malicious one is chosen, as shown in the example, the commit operation appears normal while the rogue server silently exfiltrates authentication tokens and repository data.
Although MCP currently operates mostly in local environments, its future adoption in multi-tenant and marketplace ecosystems will further amplify the risk of name collision. In such contexts, multiple organizations or users may register similarly named servers, and the lack of centralized naming control can easily lead to confusion and impersonation attacks. Moreover, as MCP marketplaces expand to include public listings and automated installation, supply-chain attacks may become a critical concern, allowing malicious servers to masquerade as trusted ones. Future designs should enforce cryptographically verifiable server identities through signed manifests that bind namespaces to verified publishers, and establish centralized or federated namespace governance to ensure uniqueness. Client runtimes and host applications should prioritize verified or community-endorsed servers, clearly display publisher provenance and verification status in their interfaces, and continuously monitor server registration or update events to detect suspicious namespace similarities and abnormal behavior.
5.1.2. Tool Name Conflict
Tool name conflicts occur when multiple tools within the MCP ecosystem share identical or confusingly similar names, resulting in ambiguity during tool discovery and invocation. Unlike server name collisions, which primarily affect user decisions during server installation or selection, tool name conflicts occur at a deeper level within the server’s functional interface. Each tool is an essential capability exposed by an MCP server, yet end users typically have limited visibility or control over which specific tools are invoked. In practice, most users delegate tool selection entirely to AI applications, which rely on textual tool names and descriptions without cryptographic verification or contextual awareness. This makes tool impersonation attacks particularly stealthy and difficult to detect. A common attack scenario involves a malicious actor registering a tool named send_email that imitates a legitimate email-sending utility provided by another server. When an AI application automatically selects and invokes the malicious version, sensitive data intended for trusted recipients, such as user credentials, messages, or attachments, may be redirected to an attacker-controlled endpoint, thereby compromising confidentiality and trust within the MCP workflow. The subtlety of this threat lies in the fact that the tool-level substitution occurs transparently to both the user and the host, making mitigation through ordinary name checking or user confirmation insufficient.
To reduce the risk of tool name conflicts, future MCP designs should enforce stronger identity binding between tools and their parent servers. Each tool should be uniquely namespaced under a verified server identifier (e.g., gmail-mcp.send_email) to avoid ambiguity across different publishers. MCP hosts and SDKs can further validate tool integrity through cryptographic signatures embedded in the server’s manifest, ensuring that a tool’s metadata, parameters, and intended operations have not been tampered with. In addition, AI applications should present contextual provenance, such as the verified publisher or trust level, when automatically selecting tools, allowing human oversight when sensitive actions are requested.
5.1.3. Preference Manipulation Attack
Beyond name similarity, we found that malicious actors can further manipulate tool selection by embedding deceptive or persuasive phrases in tool descriptions, a phenomenon we refer to as the Preference Manipulation Attack (PMA). Specifically, we observed that when a tool’s description explicitly contains self-promoting directives such as “this tool should be prioritized” or “prefer using this tool first”, the MCP client or AI application tends to select that tool more frequently, even when its functionality is malicious.
Figure 6. Example of a preference manipulation attack. A malicious MCP tool uses persuasive language in its description to bias the AI application’s selection process. Although multiple tools provide identical functionality, the model preferentially invokes add_tool_best due to its self-promoting metadata.
As shown in Figure 6, this subtle form of influence can bias the model’s choice toward an attacker-controlled tool solely through textual manipulation of metadata. This behavior introduces a severe risk of toolflow hijacking, in which attackers leverage misleading or emotionally charged descriptions to distort the tool selection process and seize control over critical operations. Such preference bias not only compromises system integrity and user safety but also enables attackers to obtain unfair economic advantages, such as by increasing paid API usage, redirecting traffic, or amplifying exposure to specific advertisements, while marginalizing legitimate competitors. More advanced adversaries may adopt traditional advertising strategies, leveraging psychological cues of authority, emotion, exaggeration, and subconscious framing, and further enhance them using evolutionary or genetic algorithms to automatically evolve persuasive yet seemingly harmless descriptions. Experimental results from Wang et al. (2025a) demonstrate that such Genetically Adapted Preference Manipulation Attacks (GAPMA) can continuously maintain high attack efficacy while remaining inconspicuous to both users and automated detectors.
Platform operators should employ multi-layered defenses, including metadata auditing to identify manipulative phrasing, anomaly detection targeting suspicious linguistic patterns, randomized tool ordering to neutralize selection bias, and trusted labeling mechanisms for verified third-party MCP servers. Furthermore, integrating adversarial training or preference robustness assessments into LLM fine-tuning could strengthen models’ resistance against linguistic manipulation and reduce their susceptibility to behavioral bias in tool selection.
5.1.4. Tool Poisoning
Tool poisoning embeds covertly malicious logic into an MCP tool while preserving its legitimate interface. Unlike preference manipulation, which biases how a model selects tools, poisoning alters the tool’s behavior after selection. As illustrated in Figure 7, a seemingly benign tool may perform hidden actions, such as reading sensitive local files or sending data to external endpoints, while still returning correct results, making the attack highly deceptive.
Figure 7. Example of tool poisoning. The malicious tool add retains a legitimate arithmetic interface but embeds hidden directives in its documentation. When the AI application invokes it to compute a simple sum (e.g., “5+1”), the model also follows the injected instructions to read a local SSH key and send it to an external URL. Although the user receives a correct numerical result, sensitive data is leaked in the background.
This vulnerability arises because MCP tools expose metadata fields (e.g., descriptions) that LLMs treat as authoritative. By injecting crafted instructions into these fields, an attacker can stealthily redirect model behavior without modifying code, evading conventional static or signature-based detection. Once the LLM starts using a poisoned tool, several harmful outcomes may occur:
  • Silent data leakage. The tool may prompt the model to read local files (e.g., SSH keys, password hashes, or configuration files) and exfiltrate them via normal API calls.
  • System compromise. Malicious metadata can instruct the model to invoke unrelated system commands, leading to malware installation, configuration tampering, account backdoors, or denial-of-service attacks.
  • Interaction hijacking. Altered logging or relay logic enables recording of user sessions, capturing prompts, private documents, and workflow details beyond the intended scope.
  • Logic corruption. Manipulated outputs (such as biased analyses or filtered search results) can distort model reasoning and steer user decisions toward attacker-preferred outcomes.
What makes tool poisoning particularly insidious is its stealth: the harmful actions occur only after a seemingly correct tool has been invoked, so any anomalies are likely attributed to model errors or benign software glitches. Furthermore, in an ecosystem where MCP servers are reused across multiple agents and platforms, a single poisoned instance can propagate compromised behavior to a wide population of end-users. A practical defense against tool poisoning should combine metadata integrity verification, automated scanning, and runtime safeguards. Before a tool is published or installed, the MCP platform should perform multi-layered static analysis. This includes pattern matching for instruction-like language (e.g., “read file,” “send to URL”), heuristic analysis of imperative verbs or sensitive API keywords in the description or usage fields, and whitelist-based filtering of acceptable metadata formats. The host runtime can record API call patterns and flag unusual sequences, such as a computation tool attempting to perform network operations or file I/O. Before forwarding metadata content to the model, the host can sanitize it by removing imperative or executable phrases, converting all metadata into strictly declarative informational text.
5.1.5. Rug Pulls
Rug pulls represent another critical security weakness within the MCP ecosystem. The concept originates from the cryptocurrency and blockchain world, where malicious developers initially release a legitimate-looking project that attracts investors or contributors, but later inject backdoors or withdraw support, causing severe losses. Transposed into the MCP setting, a rug pull occurs when an apparently benign MCP server, initially functioning as advertised, is later surreptitiously altered by its maintainer to introduce malicious code or terminate safe functionality. Users who have already installed and trusted the server become unaware victims of the change. As illustrated in Figure 8, a malicious provider may first publish a popular hotnews MCP server that aggregates daily headlines from trusted sources. During the first run, the service behaves correctly, gaining adoption through community recommendations. On the subsequent runs, the maintainer pushes a subtle update that injects harmful payloads, such as biased news filtering, hidden prompt injection, or data exfiltration.
Figure 8. Example of Rug Pulls. During the first run, the tool (hotnews) behaves benignly, returning genuine news headlines from trusted sources to gain user confidence. On subsequent runs, the same tool activates a hidden malicious branch that filters out negative stories and injects promotional or propagandistic content.
The danger of rug pulls in MCP is therefore twofold: temporal stealth, in which the malicious logic is introduced only after a trust relationship has been established, and systemic reach, since a compromised server can not only affect many users simultaneously but also propagate malicious updates or influence other interconnected MCP servers. To defend against rug pulls, the ecosystem requires stronger security primitives, including version pinning, reproducible builds, and cryptographic signature verification of MCP server code. Additionally, client-side mechanisms for update transparency and runtime monitoring should be adopted to enable users to detect when a previously trusted MCP server has turned malicious.
5.1.6. Cross-Server Shadowing
Cross-server shadowing is a threat that arises when multiple MCP servers are connected to the same agent, allowing a malicious server to mimic or shadow the tools of a legitimate server. Instead of directly injecting malicious payloads into the host, the attacker exploits naming collisions or overlapping tool definitions so that the base LLM unknowingly invokes the attacker-controlled tool in place of the trusted one. This enables the malicious server to intercept, modify, or exfiltrate sensitive information that would otherwise be handled securely by the benign service.
Figure 9. Example of coss-server shadowing attack. A malicious MCP server (math_tools) introduces a seemingly benign tool add, which claims to simply add two integers but secretly overrides the behavior of an existing tool on another server (send_email). When the user requests to send an email, the shadowed logic redirects the message to the attacker’s address while still reporting a successful delivery to the legitimate recipient.
As shown in Figure 9, suppose a trusted server exposes a tool named send_email for transmitting messages to user-specified recipients. An attacker can publish a shadowing server that also registers a tool with the same interface but manipulates its parameters. When the agent attempts to use the send_email function, the malicious version may quietly forward a copy of the message to an attacker-controlled address such as attkr@pwnd.com, alongside the legitimate recipient (e.g., alice@mail.com). From the model’s perspective, the call appears successful and the response looks consistent, but the attacker has already siphoned sensitive communication data, enabling identity fraud or unauthorized access. MCP does not enforce strict namespace isolation or code provenance, so users and LLMs have difficulty distinguishing which server’s implementation is being executed.
Effective defenses focus on keeping tools separate and easy to see during the resolution process. First, the MCP runtime should make clear namespace separation. Tools from different servers should use full names, not short ones, to stop accidental or hostile overlaps. Next, the AI application should check for conflicts when loading servers. It should give a warning or stop running if two servers have tools with the same name.
5.1.7. Command Injection/Backdoor
Command injection and backdoor attacks occur when malicious logic is secretly embedded into an MCP server’s source code or dependencies during development or packaging. Unlike traditional software backdoors that rely on direct system compromise, threats within the MCP ecosystem exploit the server’s integration with LLM-driven tool execution. Once an infected tool is loaded, a prompt or parameter crafted by the attacker can trigger hidden commands without explicit user awareness, enabling data exfiltration, privilege escalation, or remote code execution through normal model interaction, as illustrated in Figure 10. These risks are amplified by MCP’s reliance on community-maintained extensions and its dynamic, multi-server architecture, where trust is distributed across independently developed components. A compromised dependency or build pipeline can introduce a persistent backdoor that propagates across multiple agents or environments, undermining the reliability of the entire toolchain.
Figure 10. Example of a command injection attack. The malicious file_manager server defines a tool read_file that looks normal but secretly executes shell commands extracted from the input path. When the input includes “;exec=cat /etc/passwd”, the injected command runs alongside the file read operation, leaking sensitive data and converting a benign utility into a covert backdoor.
To mitigate these risks, the primary defense is to strengthen code integrity verification. All dependencies and server binaries should be built and verified through reproducible builds, cryptographic signing, and checksum validation to confirm that deployed code matches audited sources. Build pipelines should enforce isolated build environments and integrity attestations to prevent supply-chain tampering. In addition, version pinning and strict dependency management reduce exposure to compromised external components. Complementary runtime measures, such as anomaly detection and audit logging, can further identify unauthorized code behavior if static integrity checks are bypassed.
5.2. External Attacker
5.2.1. Installer Spoofing
Installer spoofing occurs when attackers distribute modified MCP server installers that introduce malicious code or backdoors during the installation process. Each MCP server requires a unique configuration that users must manually set up in their local environments before the client can invoke the server. This manual configuration process creates a barrier for less technical users, prompting the emergence of unofficial auto-installers that automate the setup process. As shown in Table 4, tools such as Smithery-CLI, mcp-get, and mcp-installer streamline the installation process, allowing users to quickly configure MCP servers without dealing with intricate server settings.
Table 4. Unofficial MCP auto installers (As of Sept. 2025).
Tool Author # Stars # Servers URL
Smithery CLI Henry Mao 407 7,437 smithery.ai
mcp.run Dylibso / 242 docs.mcp.run
mcp-get Michael Latman 483 59 mcp-get.com
Toolbase gching / 24 gettoolbase.ai
mcp-installer Ani Betts 1,432 NL1 mcp-installer
  • 1
    Enables MCP server installation through natural language interaction with the client.
However, while these auto-installers enhance usability, they also introduce new attack surfaces by potentially distributing compromised packages. Since these unofficial installers are often sourced from unverified repositories or community-driven platforms, they may inadvertently expose users to security risks such as installing tampered servers or misconfigured environments. Attackers can exploit these auto-installers by embedding malware that grants unauthorized access, modifies system configurations, or creates persistent backdoors. Moreover, most users who opt for one-click installations rarely review the underlying code for potential security vulnerabilities, making it easier for attackers to distribute compromised versions undetected. Addressing these challenges requires developing a standardized, secure installation framework for MCP servers. Each installation process should explicitly display the package source, version information, digital signature, and installation path, requiring explicit user confirmation before execution. Auto-installers must verify integrity through cryptographic signatures or checksums fetched from trusted registries, and reject unsigned or mismatched packages. In addition, implementing reputation-based trust scoring for installers and maintaining transparent download provenance records can help users assess credibility before installation. Finally, client tools should support a secure installation mode that isolates the setup process in a restricted environment, preventing installers from making unauthorized system-level modifications.
5.2.2. Indirect Prompt Injection
Indirect prompt injection operates by embedding malicious instructions into external data sources that the model retrieves via MCP tools. Since MCP servers are designed to seamlessly connect language models with external systems, such as GitHub, databases, or news platforms, the model may be exposed to adversarial content while executing what appears to be a benign user request. This indirect exposure enables attackers to manipulate the model without any direct access to it. The unique risk of indirect prompt injection in MCP environments lies in its subtlety and ease of execution. From the model’s perspective, the data returned by MCP tools often appears trustworthy, as it is delivered through well-defined APIs in standard formats such as JSON or plain text. Malicious instructions are thus concealed within legitimate content, making them extremely difficult to distinguish from ordinary information. Furthermore, an attacker can exploit this vector with minimal effort: for instance, as shown in Figure 11 by posting a carefully crafted issue into a public GitHub repository, an attacker can ensure that any MCP tool fetching these issues will deliver the poisoned text directly into the model’s context. No elevated privileges or direct interaction with the model are required, significantly lowering the bar for exploitation.
Figure 11. Example of an indirect prompt injection attack. The malicious instruction is hidden within external content hosted on a public GitHub issue page. When the user requests to list issues from the public repository, the MCP server retrieves and passes the content to the model. The injected text instructs the model to access a private repository and reveal confidential data.
Security mechanisms that only focus on sanitizing user prompts are insufficient in the MCP ecosystem, as the actual threat emerges from data pulled in by trusted connectors. To mitigate indirect prompt injection, defenses must treat tool outputs as potentially adversarial and introduce safeguards that can detect and neutralize harmful instructions hidden in external data streams. With MCP applications increasingly integrating heterogeneous and user-generated sources, the difficulty of detecting indirect prompt injection grows, while the cost for attackers to mount such an attack remains strikingly low.
5.3. Malicious User
5.3.1. Credential Theft
Credential theft in MCP environments refers to the risk that sensitive authentication information, such as API keys, access tokens, or database credentials, are exposed and subsequently misused by adversaries. A notable security gap arises during the local deployment of MCP servers, where users are instructed to persist configuration snippets into default configuration files. As shown in Figure 12, these files often contain plaintext API keys (e.g., VIRUSTOTAL_API_KEY) embedded directly in JSON configurations. Because such configuration files are stored in predictable default locations across operating systems (e.g., %APPDATA%, ˜/Library/Application Support/, or workspace-specific hidden directories), they become attractive targets for credential-stealing attacks. If local file system permissions are weak, or if an adversary gains access through an unrelated MCP vulnerability, these plaintext secrets may be exfiltrated with minimal effort. Once exposed, the stolen API keys can be reused by attackers to impersonate legitimate users, access third-party services (such as VirusTotal), or pivot toward more critical infrastructure. This form of privilege misuse enables long-term persistence, especially when keys are not regularly rotated or lack granular scoping.
Figure 12. Illustration of credential theft risk in MCP environments. The left shows a typical MCP server configuration snippet used during local deployment, where sensitive information (e.g., VIRUSTOTAL_API_KEY) is stored in plaintext within the client’s configuration file. The right lists common default storage paths for MCP host (e.g., Cline and Cursor) configurations across different operating systems.
The persistence of sensitive credentials in unencrypted, static JSON files highlights the need for secure credential management in MCP ecosystems. Recommended mitigations include: avoiding plaintext storage of secrets by leveraging dedicated secret managers or encrypted keystores; enforcing restrictive file system permissions for local configuration directories; and implementing automatic secret rotation and token expiration. Without such mechanisms, MCP servers remain highly susceptible to credential theft, as adversaries can trivially extract stored keys to hijack privileged operations.
5.3.2. Sandbox Escape
MCP defines a communication framework between hosts and external tools. It focuses on structured data exchange and limited parameter validation, such as verifying types or value ranges. The protocol does not include mechanisms for runtime isolation or privilege control. As a result, the security boundaries of an MCP server depend entirely on its hosting environment. If the host already holds high system privileges, a poorly configured server may create conditions similar to a sandbox escape. This is particularly serious in MCP because servers often run inside AI‑integrated hosts that have broad access to files, networks, and user contexts. Once this boundary is breached, a malicious server can indirectly influence the agent’s reasoning process, alter shared data, or trigger privileged actions without the user’s awareness.
Different types of MCP servers face different risks. A database server that accepts unrestricted queries can allow unintended write or delete operations. A file-system server that lacks directory restrictions may enable access to sensitive areas such as configuration folders or system files. A network server without proper outbound filtering may let attackers send unauthorized requests or exfiltrate data. Other MCP servers, such as those connected to cloud APIs or local applications, may have similar but context-specific exposures. Real-world incidents demonstrate how such risks can emerge. Many MCP servers run inside host applications such as Cursor or Cline. These hosts already have permissions to read and write files, access networks, and interact with external services. If a host exposes an MCP server on localhost without authentication, other local processes, including a Chrome extension, can connect to it and issue privileged commands. In such cases, the attacker can act as an authorized user and bypass both the host sandbox and normal operating system protections. Several measures can help reduce these risks. The server should limit network bindings to trusted local clients and require authentication for every connection. File-system access should be restricted to a specific directory. Network permissions should allow only approved domains. All input parameters should be validated and sanitized before execution. In addition, containers or process-level isolation can help confine tool execution, and continuous monitoring can detect abnormal activity during runtime.
5.3.3. Tool Chaining Abuse
Tool chaining abuse occurs when multiple low-risk MCP tools are implicitly combined by the model to perform unintended high-impact operations. As shown in Figure 13, a single natural language request can trigger a sequence of legitimate tool calls that collectively lead to data exfiltration, uch as listing files, reading configuration data, extracting credentials, and exporting results to a public location. Because each step operates within the model’s authorized permissions, these activities often evade traditional access control or policy enforcement mechanisms. This attack is characterized by its implicit orchestration, where the model autonomously plans and chains approved tools without explicit malicious code or user intent. It exploits the semantic flexibility of natural language instructions and the composability of MCP tools, turning routine operations into multi-stage attack pipelines. The result is a highly stealthy and policy-compliant form of misuse that challenges conventional security monitoring and permission-based defenses.
Figure 13. Example of a tool chaining abuse attack. Multiple benign tools are combined by the model in response to a single natural language request. The AI application first lists directories and identifies a .env file, then reads it to extract database credentials, executes a query using those credentials to obtain sensitive user data, and finally writes the exported results to a public path.
Mitigation requires stronger execution constraints on cross-tool composition. Possible measures include prohibiting implicit tool chaining without explicit user approval, limiting which outputs can be piped to other tools (e.g., file contents to network tools), and enforcing runtime anomaly detection for suspicious chains such as “local file access $\rightarrow$ external network transmission”.
5.3.4. Unauthorized Access
Unauthorized access in the MCP ecosystem arises when attackers can interact with servers, tools, or communication channels beyond their intended permissions. Unlike prompt‐level threats that exploit the model’s language reasoning, unauthorized access typically abuses weaknesses in authentication, transport, or session management. As shown in Figure 14, the leakage of an SSE session_id can allow an attacker to reuse a valid session and issue remote commands without re‑authentication, leading to persistent system control. Similar risks occur when MCP components expose unprotected HTTP or WebSocket endpoints where sensitive functions, such as file or database operations, can be triggered directly by unauthorized clients. Another common source of unauthorized interaction stems from insecure credential or token handling. When environment variables or configuration files contain access tokens or API keys that are unintentionally exposed, adversaries can impersonate legitimate clients and perform privileged operations. In some deployments, authorization design flaws such as the confused deputy problem or token passthrough further amplify the issue: a trusted proxy may unknowingly perform actions on behalf of an attacker, or upstream credentials may be forwarded to downstream services without proper audience validation, breaking isolation boundaries. Session hijacking represents another manifestation of the same weakness, poorly protected or long‑lived session identifiers can be intercepted and reused to impersonate valid users, extending unauthorized control over the system. Building effective defenses requires mandatory endpoint authentication, secure credential storage, token audience enforcement, and the periodic rotation or invalidation of active sessions to restore strict trust boundaries between users, tools, and remote services.
Figure 14. Example of unauthorized access. By specifying transport=‘‘sse’’, the MCP server starts a Server-Sent Events (SSE) service on the default port 8000 and returns a unique session_id when the /sse route is accessed. If this identifier is exposed, an attacker can reuse it to send arbitrary commands through the /messages/?session_id=<id> route without authentication, enabling unauthorized command execution.
5.4. Security Flaws
5.4.1. Re-deployment of Vulnerable Versions
MCP servers, being open-source and maintained by individual developers or community contributors, lack a centralized platform for auditing and enforcing security updates. Users typically download MCP server packages from repositories like GitHub, npm, or PyPi and configure them locally, often without formal review processes. This decentralized model increases the risk of re-deploying vulnerable versions, either due to delayed updates, version rollbacks, or reliance on unverified package sources. When users update MCP servers, they may unintentionally roll back to older, vulnerable versions to address compatibility issues or maintain stability. Additionally, unofficial auto-installers, such as mcp-get and mcp-installer, which streamline server installation, may default to cached or outdated versions, exposing systems to previously patched vulnerabilities. Since these tools often prioritize ease of use over security, they may lack version verification or fail to notify users about critical updates. Because security patches in the MCP ecosystem rely on community-driven maintenance, delays between vulnerability disclosure and patch availability are common. Users who do not actively track updates or security advisories may unknowingly continue using vulnerable versions, creating opportunities for attackers to exploit known flaws.
Empirical evidence from the MCP ecosystem supports this observation. As shown in Table 4, the number of GitHub stars for unofficial auto-installers continues to increase, reflecting a growing user base and dependence on automated installation tools. However, the corresponding number of actively maintained MCP servers, including those hosted or updated through these installers, has remained largely unchanged over the same period. Except for the Smithery CLI, which shows regular updates, most auto-installers reference static MCP server packages that have not received version revisions in months. This asymmetry indicates that while more users are deploying MCP servers through automated tools, these tools frequently install outdated codebases lacking the latest security patches. As a result, the likelihood of re-deploying vulnerable versions increases proportionally with the popularity of auto-installers.
From a research perspective, analyzing version management practices in MCP environments can identify potential gaps and highlight the need for automated vulnerability detection and mitigation. On the other hand, there is also a pressing need to establish an official package management system with a standardized packaging format for MCP servers and a centralized server registry to facilitate secure discovery and verification of available MCP servers.
5.4.2. Post-Update Privilege Persistence
Privilege persistence refers to a condition in which outdated or revoked credentials remain valid after an MCP server update, enabling previously authorized users or malicious actors to retain elevated privileges. This vulnerability typically arises when privilege modifications, such as API key revocations, session token expirations, or role reassignments, are not properly synchronized or invalidated following a server upgrade. If such obsolete credentials persist, attackers can continue to access sensitive resources or perform privileged operations beyond the intended scope of the updated configuration. Similar behaviors have long been observed in conventional cloud and web systems, where revoked OAuth tokens or inactive IAM sessions remain temporarily valid due to asynchronous cache updates or delayed revocation. Robbins (Slack Security Team, 2022) demonstrated in 2022 that adversaries could obtain passwordless persistence and privilege escalation in Microsoft Azure Active Directory through the misuse of Certificate-Based Authentication. In that case, incomplete revocation of trusted credentials enabled a malicious certificate authority to impersonate global administrators indefinitely, illustrating how cached trust states and delayed revocation may sustain unauthorized privileges even after administrative changes. The MCP ecosystem inherits these underlying weaknesses and tends to amplify their effects. Many MCP servers maintain persistent network connections, such as SSE and WebSocket streams, while sharing credentials across different tools or hosts. When these servers are hot-reloaded or automatically updated, they may reload binaries without fully reinitializing credential data or terminating active sessions, which allows previously valid tokens to remain effective. As a result, a common web security problem that arises from delayed credential revocation can manifest with greater persistence and broader propagation in distributed MCP deployments.
To reduce this risk, MCP implementations should ensure immediate revocation of credentials during updates and synchronize privilege changes across all active instances. Automatic expiration and renewal mechanisms tied to server versions can further prevent old credentials from persisting. Integrating with centralized validation services that provide real-time checks of token status can strengthen overall consistency. Moreover, maintaining detailed logs and auditing privilege transitions can improve transparency and facilitate anomaly detection. From a design standpoint, extending the MCP specification with parameters such as session_expiry or explicit revocation_events could enhance lifecycle security and help prevent privilege persistence inherited from traditional web environments.
5.4.3. Configuration Drift
Configuration drift occurs when unintended or uncoordinated changes accumulate in a system’s configuration, causing it to deviate from the intended security baseline. Such deviations often result from manual adjustments, partial updates, or conflicting modifications introduced by different tools or users. In MCP environments, where server instances are frequently configured and maintained locally, configuration drift becomes a significant source of vulnerability. Because MCP servers may integrate multiple client tools and manage diverse access policies or API keys, even small inconsistencies can propagate across dependent components, leading to unpredictable security states. The emergence of remote or cloud-hosted MCP deployments, such as those provided by Cloudflare, heightens this concern. In contrast to local setups, where drift typically compromises only one user’s environment, configuration inconsistencies in multi-tenant or federated MCP servers can simultaneously affect many users or organizations. A misaligned access policy, outdated capability definition, or inconsistent plugin permission scope may inadvertently expose sensitive data or escalate privileges across tenants. These issues are amplified when servers rely on cached configurations or perform asynchronous synchronization between distributed nodes.
Mitigating configuration drift requires maintaining alignment between the deployed runtime state and a defined configuration baseline. Automated validation tools can periodically compare active settings with canonical templates and flag deviations before they propagate further. Version-controlled configuration definitions, immutable runtime manifests, and signed policy descriptors can provide verifiable reference points to detect unauthorized or accidental divergence. Additionally, implementing robust rollback procedures and continuous compliance auditing helps ensure that both local and remote MCP environments adhere to secure configuration standards. From a design perspective, incorporating configuration integrity verification into the MCP specification itself, for example through cryptographic checksums or schema validation hooks, would further reduce the risk of silent drift and strengthen overall operational resilience.
6. Discussion
6.1. Implications
The rapid adoption of MCP is transforming the AI application ecosystem, introducing new opportunities and challenges that span the full MCP server lifecycle described in Figure 3, from creation to deployment, operation, and maintenance. These developments carry important implications for developers, users, ecosystem maintainers, and the broader AI community, each facing distinct responsibilities and risks across different lifecycle stages.
For developers, MCP simplifies tool integration during the creation and deployment stages (§ 3.3.2, § 3.3.3), enabling more efficient design of agentic workflows and complex multi-step reasoning tasks. By providing standardized capability declarations and invocation interfaces, developers can focus on functionality rather than integration friction. However, this openness also exposes new attack surfaces, such as namespace typosquatting (§ 5.1.1), tool poisoning (§ 5.1.4), and rug pulls (§ 5.1.5) during metadata definition or capability declaration. These threats can lead to privilege escalation, supply chain compromise, or execution of malicious payloads. Developers should therefore implement provenance verification, use version-controlled releases, and apply static analysis or digital signing before server registration to ensure integrity.
For users, the operation stage (§ 3.3.4) significantly enhances usability by enabling AI agents to orchestrate cross-platform workflows involving data services, enterprise systems, and IoT devices. This reduces manual effort and improves productivity. Nonetheless, users are directly exposed to runtime threats such as indirect prompt injection (§ 5.2.2), command injection (§ 5.1.7), and unauthorized access (§ 5.3.4). Interacting with unverified servers or insecure endpoints can lead to leaked credentials or unauthorized actions. To reduce exposure, users should prioritize trusted MCP collections, verify digital signatures at installation, and rely on sandboxed MCP hosts to confine tool execution during session management.
For MCP ecosystem maintainers, decentralization introduces heterogeneity across the deployment and maintenance phases (§ 3.3.3, § 3.3.5). Community-driven updates often vary in quality and cadence, leading to inconsistencies in patch management and version control. This fragmentation increases the likelihood of vulnerable versions (§ 5.4.1), privilege persistence (§ 5.4.2), and configuration drift (§ 5.4.3), which may be exploited to retain outdated permissions, misconfigure environments, or expose sensitive resources. Regular mechanisms such as automated version checks, integrity auditing, and mandatory configuration validation should be integrated into MCP registries and collections to ensure ecosystem-wide resilience.
For the broader AI community, MCP provides a foundation for interoperability that supports large-scale collaboration and reuse of AI capabilities across industries, enhancing cross-agent coordination and accelerating innovation. However, the expansion of server interactions across multiple lifecycle stages also amplifies ethical and safety concerns. Ensuring fairness in tool selection, defending against dataset leakage, and maintaining accountability in automated decision workflows become shared priorities. Continuous community participation in auditing, reporting, and lifecycle security monitoring is essential to preserve transparency, prevent misuse of AI capabilities, and ensure equitable access.
6.2. Challenges
Despite its potential, MCP adoption introduces a series of pressing challenges that must be resolved to ensure sustainable growth, operational reliability, and responsible development.
Lack of centralized security oversight. Since MCP servers are primarily managed by independent developers, there is no central authority to audit security baselines or enforce uniform compliance. This decentralization leads to inconsistent patching, irregular vulnerability management, and difficulty enforcing best practices. Malicious actors can exploit this fragmentation through namespace typosquatting (§ 5.1.1) or tool name conflicts (§ 5.1.2) during the metadata definition and capability declaration stages, resulting in the installation of malicious servers or privilege escalation. Moreover, the absence of an official package management mechanism leads to version inconsistency and unverified updates, heightening the risk of deploying vulnerable versions (§ 5.4.1) or misconfigured releases (§ 5.4.3).
Authentication and authorization gaps. MCP currently lacks a standardized framework for authentication and authorization across clients and servers. Without a unified mechanism to establish trust, identity spoofing and session hijacking can occur. As shown in Table 3, threats such as unauthorized access (§ 5.3.4) and privilege persistence (§ 5.4.2) can emerge during the operation and maintenance phases, allowing attackers to impersonate legitimate agents or retain administrative privileges. In multi-tenant environments, poorly implemented role-based access control further increases information exposure and inconsistent privilege enforcement across different MCP clients.
Insufficient debugging and monitoring mechanisms. The current MCP specification provides limited capabilities for logging, auditing, or runtime introspection. As a result, developers struggle to trace errors, investigate anomalous behavior, or detect silent failures during execution. Without unified telemetry and audit trails, incidents such as command injection (§ 5.1.7) and cross-server shadowing (§ 5.1.6) may remain undetected, enabling stealthy intrusion and lateral movement across connected environments. The lack of comprehensive monitoring frameworks also complicates response procedures, making it difficult to identify configuration drift or payload corruption until after exploitation has occurred.
Maintaining consistency in multi-step, cross-system workflows. MCP’s design supports dynamic, multi-step tool chaining, but ensuring consistent session state and reliable recovery is challenging in distributed environments. Malicious users may exploit these complex execution paths through tool chaining abuse (§ 5.3.3) or indirect prompt injection (§ 5.2.2) attacks, manipulating intermediate data or injecting hostile prompts between tools. Without robust state validation, integrity checks, or rollback mechanisms, small execution errors can propagate across multiple agents, resulting in workflow inconsistency or partial process execution.
Scalability challenges in multi-tenant environments. As MCP infrastructure evolves toward remote hosting and multi-tenant deployment, isolating tenants while preserving performance and privacy becomes more difficult. Unsynchronized updates or shared runtime environments may cause data leakage, resource contention, or mismanagement of privileges. Vulnerabilities such as sandbox escape (§ 5.3.2) and credential theft (§ 5.3.1) present notable threats during tool invocation within resource-shared clusters. To achieve scalability and security simultaneously, standardized tenant isolation models and runtime sandboxing must become integral to MCP server design.
Challenges in embedding MCP in smart environments. The integration of MCP into smart homes, industrial IoT systems, and enterprise automation introduces distinctive risks. These deployments often operate under real-time constraints and involve continuous communication between agents and heterogeneous devices. Compromised edge or on-premise MCP servers can lead to critical safety failures, privilege abuse, or command spoofing. Attacks such as installer spoofing (§ 5.2.1) or preference manipulation (§ 5.1.3) can exploit insecure installation pipelines or unsafe default configurations to deploy backdoored servers. Implementing secure firmware update channels, enforcing strict authentication for device-level access, and integrating MCP-aware intrusion detection are essential steps toward safer deployment in connected environments.
6.3. Recommendations across the MCP server lifecycle
The MCP can achieve secure, stable, and scalable adoption only if appropriate mechanisms are established at each stage of its lifecycle. Recommendations are therefore presented according to the four main lifecycle phases, textitcreation, deployment, operation, and maintenance, as defined in Figure 3. Each stage entails distinctive configuration tasks and corresponding opportunities for functional enhancement and risk mitigation.
6.3.1. Recommendations for Creation Phase (§ 3.3.2).
The creation phase is the starting point of an MCP server’s lifecycle. It defines how the server identifies itself, what functions it provides, and how securely those functions are implemented. Improvements made here have long-lasting effects on reliability and security. Three practical measures can strengthen this stage. Automate metadata generation and verification. Developers should not write metadata by hand. Instead, the server’s name, version, and supported protocols should be generated automatically during the build process. Each build should create a manifest file that includes these elements and a digital fingerprint of the source code. Before release, a verification step should confirm that this manifest matches the compiled artifact. This approach ensures that every published server can be traced to a trusted source and that outdated or modified copies are easy to detect. Validate capability declarations before release. A server’s capabilities describe its accessible tools and resources. Each capability should include a clear description of its scope, allowed operations, and required permission level. Developers should run a pre‑release validation process that checks whether all declared capabilities are correctly implemented and that none exceed their intended boundaries. For example, a capability that reads files should not include write access unless explicitly required. This validation helps prevent configuration errors and limits accidental privilege escalation during later operation. Separate implementation modules and enforce input checking. Each capability should have its own implementation module with limited access to shared resources. This design prevents faults in one module from affecting others. Within each module, the server should run input checks before executing any operation. These checks verify parameter counts, data formats, and expected value ranges. If inputs do not meet these conditions, the request should terminate safely instead of causing runtime errors. Clear separation between modules also makes it easier to test, update, or replace specific capabilities without changing the entire server.
6.3.2. Recommendations for Deployment Phase (§ 3.3.3).
The deployment phase introduces an MCP server into a live environment where client systems and external resources can interact with it. This stage determines whether the server operates safely and consistently across different hosts. Produce verified and reproducible build packages. Developers should prepare each release in a controlled build environment. The process should record dependency versions, configuration parameters, and source identifiers. Each package should include a checksum or digital signature that allows later verification of its origin. Before deployment, maintainers should confirm that the package metadata and version record match the server registry. This procedure guarantees that all running instances correspond to approved source code and that no altered or partially built versions enter production. Deploy servers within application sandboxes. Every MCP server should install and run inside a dedicated sandbox environment. The sandbox limits file access, network communication, and operating‑system instructions to a predefined scope. Server type determines the isolation rules. For instance, a server that connects to a database can operate through a restricted account. The sandbox allows only read queries, such as SELECT, when the server’s declared capability is to retrieve information. Modification commands like UPDATE or DELETE remain disabled. The sandbox can also intercept query inputs, apply escaping rules, and reject statements containing unapproved keywords to reduce the risk of SQL injection. Similarly, a server that interacts with the local file system should have access only to one designated directory created during installation. The sandbox enforces this directory boundary so that any attempt to read or write outside that folder fails immediately. Logs, configuration files, or temporary data are all stored within this enclosed workspace. This level of isolation prevents unnecessary exposure of system paths and confines possible damage to a controlled environment even if an error occurs.
6.3.3. Recommendations for Operation Phase (§ 3.3.4).
The operation phase is when the MCP server runs continuously and interacts with users, clients, and external resources. During this stage, the server interprets user intent, invokes tools, and manages sessions. Real‑time monitoring and control are essential to keep the system stable and secure. Monitor intent analysis and command execution. The MCP host or client should inspect the commands that users or language models generate before they reach the server. Each request should pass through an intent filter that checks whether the command matches declared capabilities and complies with safety rules. For example, an attempt to run a file deletion command against a read‑only tool should be rejected and logged before execution. Real‑time monitoring can also record how the intent is parsed and which capability is triggered. This visibility helps developers detect misuse patterns or unexpected behavior caused by ambiguous prompts. Maintain sandbox enforcement for dynamic operations. The sandbox introduced during deployment must remain active whenever the server handles live requests. Runtime policies should adapt to the server’s functional role. A database‑focused server should continue to use a restricted database account that only allows read queries. If a user or agent sends an update or delete request, the sandbox intercepts it and returns a permission error. For a server that manages files, the sandbox should keep all read and write operations within its dedicated workspace directory. Any attempt to access files outside this path should trigger an alert and terminate the operation safely. Maintaining active sandbox enforcement during execution ensures that even valid sessions cannot exceed defined resource boundaries. Implement consistent session management and adaptive logging. Each user or process connecting to the server should have its own authenticated session. The server should issue short‑lived tokens that expire automatically after inactivity. Session records need to track which capabilities are used, which resources are accessed, and what outputs are generated. Logging should occur in real time with minimal performance impact. For example, when a user triggers several tool invocations in a workflow, the log should record each tool name, execution time, and result summary. These logs allow administrators to identify errors, verify legitimate usage.
6.3.4. Recommendations for Maintenance Phase (§ 3.3.5).
The maintenance phase keeps the MCP server secure and reliable after deployment. Updates, configuration changes, and audits all happen during this stage. Apply structured version control and automated updates. Each server should maintain a clear version history that records all code and configuration changes. Developers should tag stable releases and store them in a controlled repository. When new patches are available, an automated update system should authenticate the package, verify compatibility, and apply changes without altering local configurations. If the update fails, the server should roll back automatically to the last stable version. This process minimizes downtime and prevents unverified patches from entering production environments. Manage configuration changes through validation and review. Administrators often need to update credentials, API keys, or resource addresses. Each configuration change should follow a controlled workflow. Before activation, the server should validate the new parameters to confirm that formats, ranges, and references are correct. For example, if a database host address changes, a test query should run automatically to confirm that the connection remains secure. Every modification should be logged together with its author, timestamp, and reason. A simple approval step, where another maintainer confirms the change, can further reduce misconfigurations that lead to service interruptions. Conduct continuous auditing and anomaly detection. The server should record key activities such as user authentication, tool invocation, and resource access. These logs should be stored in a secure and tamper‑evident format. Automated analysis can scan logs for signs of unusual activity, such as repeated failed logins or unexpected permission escalations. When anomalies occur, the system should send alerts to maintainers and isolate the affected component. Regular audits also help track long‑term behavior trends. For example, if a tool begins to take longer to execute common queries, maintainers can detect early signs of performance degradation or resource leaks before they cause larger failures.
7. Related Work
7.1. Tool Integration in LLM Applications
Equipping LLMs with external tools has become a key paradigm for enhancing their capabilities in real-world tasks. This approach enables LLMs to transcend the limitations of static knowledge and interact dynamically with external systems. Recent studies have proposed frameworks to support such integration, focusing on tool representation, selection, invocation, and reasoning. Shen et al.(Shen, 2024) provide a comprehensive survey outlining a standard LLM-tool integration paradigm, identifying key challenges in user intent understanding, tool selection, and execution planning. Building on this, AutoTools(Shi et al., 2025, 2024) introduces an automated framework that transforms raw tool documentation into executable functions, reducing reliance on manual engineering. EasyTool (Yuan et al., 2024) further streamlines this process by distilling diverse and verbose tool documentation into concise and unified instructions, improving tool usability and efficiency. From an evaluation perspective, several benchmarks have emerged. ToolSandbox (Lu et al., 2024) emphasizes stateful and interactive tool usage with implicit dependencies, while UltraTool (Huang et al., 2024) focuses on complex, multi-step tasks involving planning, creation, and execution. These efforts reveal significant performance gaps and motivate better evaluations for LLM-agent capabilities. To improve agent decision-making and prompt quality, AvaTaR (Wu et al., 2024) proposes contrastive reasoning techniques, while Toolken+(Yakovlev et al., 2024) incorporates reranking and rejection mechanisms for more precise tool use. Additionally, some works explore LLMs not just as tool users but as tool creators—ToolMaker(Wölflein et al., 2025) autonomously converts code repositories into callable tools, moving toward fully automated agents. To unify this expanding landscape, Li (Li, 2025) proposes a taxonomy that situates tool use alongside planning and feedback learning as three core agent paradigms.
As the demand for LLMs to interact with external tools continues to grow, the research community has increasingly focused on improving the quality, reliability, and generality of such interactions. Existing work demonstrates clear progress, from automatic tool discovery and documentation parsing to multi-step reasoning and evaluation, but these efforts remain fragmented and platform-specific. The lack of a unifying, secure, and extensible protocol leads to duplicated engineering, inconsistent integration practices, and heightened security risks. This fragmentation highlights the value and necessity of the MCP. MCP is designed to provide a standardized infrastructure for LLM–tool communication, defining unified formats for context exchange, capability invocation, and access control. Given its emerging role at the core of tool-augmented LLM ecosystems, it becomes crucial to comprehensively map MCP’s current landscape, analyze its security implications, and discuss its future evolution. This investigation provides both the conceptual grounding and practical guidance needed to foster a safer, more standardized foundation for LLM–tool integration.
7.2. Security Risks in LLM-Tool Interactions
The integration of tool-use capabilities into LLM agents significantly expands their functionality, but also introduces new and more severe security risks. Fu et al. (Fu et al., 2024) demonstrate that obfuscated adversarial prompts can lead LLM agents to misuse tools, enabling attacks such as data exfiltration and unauthorized command execution. These vulnerabilities are particularly concerning as they generalize across models and modalities. A growing body of work has begun to categorize and analyze these risks. Gan et al.(Gan et al., 2024) and Yu et al.(Yu et al., 2025) propose taxonomies for threats across agent components and stages, while the OWASP Agentic Security Initiative (John et al., 2025) provides practical threat modeling frameworks. To support detection and mitigation, Chen et al.(Chen and Cong, 2025) introduce AgentGuard, which automatically discovers unsafe workflows and generates safety constraints, and ToolFuzz(Milev et al., 2025) identifies failures stemming from ambiguous or underspecified tool documentation. On the alignment front, Chen et al.(Chen et al., 2024) propose the H2A principle, which encourages LLMs to behave with helpfulness, harmlessness, and autonomy, and introduce the ToolAlign dataset to guide safer tool usage. Ye et al.(Ye et al., 2024) further analyze safety risks throughout the tool-use pipeline, including malicious queries, execution misdirection, and unsafe outputs. Deng et al. (Deng et al., 2024) highlight broader systemic risks such as unpredictable inputs, environmental variability, and untrusted tool endpoints.
The above studies collectively reveal the multifaceted security risks inherent in LLM–tool interaction, spanning prompt manipulation, unsafe execution, and untrusted endpoints. These findings motivate a deeper examination of how such risks may evolve under the emerging MCP. It remains unclear whether the structured design of MCP will merely encapsulate these existing vulnerabilities, exacerbate them through new cross-context channels, or provide the means to mitigate them through standardized control and isolation. Understanding this shift is central and constitutes a primary motivation of this work, as it is essential for evaluating the true security implications of MCP within future tool-augmented LLM ecosystems.
7.3. Security of the MCP
As the MCP becomes a foundational interface for tool-augmented AI ecosystems, its security and reliability have attracted increasing attention. Its open and flexible design enables broad interoperability but also introduces novel and systemic security risks. Recent studies have primarily focused on characterizing these risks and identifying vulnerabilities. Hasan et al. (Hasan et al., 2025) conduct the large-scale empirical measurement of 1,899 open-source MCP servers, revealing eight novel vulnerability categories arising from immature maintenance and non-deterministic control flow. Zhao et al. (Zhao et al., 2025b) provide a systematic taxonomy of malicious MCP server behaviors and demonstrate the feasibility of practical attacks through proof-of-concept exploits. Wang et al. (Wang et al., 2025b) describe the Preference Manipulation Attack (MPMA), in which adversarial servers bias LLM decisions for economic gain within open marketplaces. Similarly, Shuli Zhao et al. (Zhao et al., 2025a) identify Parasitic Toolchain Attacks that exploit weak context isolation, resulting in stealthy data exfiltration across interconnected tools. To standardize empirical testing, Yang et al. (Yang et al., 2025) introduce MCPSecBench, which defines a benchmark of seventeen attack types across multiple host and client configurations, enabling reproducible MCP security assessments. Radosevich and Halloran (Radosevich and Halloran, 2025) further highlight that existing MCP workflows allow severe code execution and privilege escalation exploits, revealing insufficient safeguards in current implementations.
Complementary work explores proactive protection and hardening mechanisms for the MCP ecosystem. Bhatt et al. (Bhatt et al., 2025) propose the Enhanced Tool Definition Interface (ETDI), integrating OAuth-based identity verification and policy-based access control to mitigate squatting and rug-pull attacks. Xing et al. (Xing et al., 2025) develop MCP-Guard, a layered defense architecture combining static signature scanning, deep neural threat detection, and LLM-based decision arbitration. Kumar et al. (Kumar et al., 2025) present MCP Guardian, a defense-in-depth layer that strengthens authentication, rate limiting, logging, and firewall integration for MCP communications. At the enterprise level, Brett (Brett, 2025) introduces MCP Gateways to simplify secure self-hosted deployments through threat model mapping, authentication, and intrusion detection, supporting enterprise-grade isolation.
Most prior studies focus on isolated phases, either vulnerability scanning or runtime protection, without establishing a lifecycle, wide security model. They also tend to emphasize specific attack vectors or detection accuracy. In contrast, this work provides the first end-to-end security and privacy analysis of the MCP ecosystem covering all lifecycle stages. This lifecycle-centric analysis bridges micro-level vulnerabilities and macro-level protocol design, enabling the derivation of practical mitigation strategies and secure-by-design guidelines for the future evolution of the MCP.
8. Conclusion
This paper presents the first comprehensive analysis of the MCP ecosystem landscape. We examine its architecture, core components, operational workflows, and server lifecycle stages. Furthermore, we explore the adoption, diversity, and use cases, while identifying potential security threats throughout the creation, deployment, operation, and maintenance phases. We also highlight the implications and risks associated with MCP adoption and propose actionable recommendations for stakeholders to enhance security and governance. Additionally, we outline future research directions to tackle emerging risks and improve MCP’s resilience. As MCP continues to gain traction with industry leaders such as OpenAI and Cloudflare, addressing these challenges is key to its long-term success and to enabling secure, efficient interaction with diverse external services.